OpenStudy (laddiusmaximus):

find the derviative of y=(ln(10x))^11

OpenStudy (agreene):

\[\frac{d}{dx}=\frac{11\ln^{10}(10x)}{x}\] chain rule it up.

OpenStudy (slaaibak):

First apply chain rule: \[y' = 11(\ln(10x))^{10} * {d \over dx}(\ln(10x)) \] d/dx { ln(10x) }= x Therefore \[y' = {11(\ln(10x))^{10} \over x}\]

OpenStudy (rulnick):

11 ln(10x) ^10 * (1/(10x)) * 10 = (11/x) (ln(10x))^10, which looks like I agree with slaaibak.

OpenStudy (agreene):

we all have the same answer, seems like it's probably true.

OpenStudy (rulnick):

Ah, yes, thank you agreene, rereading your TeX I agree :)

OpenStudy (laddiusmaximus):

that isnt right

OpenStudy (rulnick):

Perhaps you meant y=ln ((10x)^11) ? The answer for that one would be y = 11/x.